OpenStudy (anonymous):

In a family with eight children, what is the probability that exactly six are boys?

OpenStudy (anonymous):

@Hope_nicole

OpenStudy (anonymous):

set it up as a fraction....6/8 if it needs to be in a percent then put it in percent form if not just reduce the fraction.

OpenStudy (anonymous):

exactly 6, need the binomial for this one

OpenStudy (anonymous):

\[\binom{8}{6}(\frac{1}{2})^8\]

OpenStudy (anonymous):

assuming of course that boys and girls are equally likely

OpenStudy (anonymous):

\(\binom{8}{6}=\frac{8\times 7}{2}=4\times 7=28\) and \(\frac{1}{2^8}=\frac{1}{256}\) so you get \[\frac{28}{256}\] reduce if you like